# DEL-Ranking: Ranking-Correction Denoising Framework for Elucidating Molecular Affinities in DNA-Encoded Libraries

### v1

@@ A.2 Gradient Analysis
A.3 Hyperparameter Setting
@@
ARDC weight 1 0.1 0.1
PSR weight 1.00e111.00E111.001011start_ARG 1.00 end_ARG start_ARG  end_ARG start_ARG e start_ARG 11 end_ARG end_ARG 1.00e91.00E91.00109start_ARG 1.00 end_ARG start_ARG  end_ARG start_ARG e start_ARG 9 end_ARG end_ARG 1.00e101.00E101.001010start_ARG 1.00 end_ARG start_ARG  end_ARG start_ARG e start_ARG 10 end_ARG end_ARG 1.00e81.00E81.00108start_ARG 1.00 end_ARG start_ARG  end_ARG start_ARG e start_ARG 8 end_ARG end_ARG 1.00e81.00E81.00108start_ARG 1.00 end_ARG start_ARG  end_ARG start_ARG e start_ARG 8 end_ARG end_ARG
LGR weight 1.0 0.1 0.1 2.0 1.0
Temperature 0.8 0.3 0.2 0.9 0.2
Contrast weight 1.00e31.00E-31.0010-3start_ARG 1.00 end_ARG start_ARG  end_ARG start_ARG e start_ARG - 3 end_ARG end_ARG 1.00e31.00E-31.0010-3start_ARG 1.00 end_ARG start_ARG  end_ARG start_ARG e start_ARG - 3 end_ARG end_ARG 1.00e31.00E-31.0010-3start_ARG 1.00 end_ARG start_ARG  end_ARG start_ARG e start_ARG - 3 end_ARG end_ARG 1.00e41.00E-41.0010-4start_ARG 1.00 end_ARG start_ARG  end_ARG start_ARG e start_ARG - 4 end_ARG end_ARG 1.00e31.00E-31.0010-3start_ARG 1.00 end_ARG start_ARG  end_ARG start_ARG e start_ARG - 3 end_ARG end_ARG
Margin 1.00e31.00E-31.0010-3start_ARG 1.00 end_ARG start_ARG  end_ARG start_ARG e start_ARG - 3 end_ARG end_ARG 1.00e31.00E-31.0010-3start_ARG 1.00 end_ARG start_ARG  end_ARG start_ARG e start_ARG - 3 end_ARG end_ARG 1.00e31.00E-31.0010-3start_ARG 1.00 end_ARG start_ARG  end_ARG start_ARG e start_ARG - 3 end_ARG end_ARG 1.00e31.00E-31.0010-3start_ARG 1.00 end_ARG start_ARG  end_ARG start_ARG e start_ARG - 3 end_ARG end_ARG 1.00e31.00E-31.0010-3start_ARG 1.00 end_ARG start_ARG  end_ARG start_ARG e start_ARG - 3 end_ARG end_ARG
A.4 Visualization of Top-50 Selection
The visualization of the Top-50 DEL-Ranking results based on predicted target counts on five datasets are shown in Figure Figures 4, 5, 6, 8 and 7. We highlight the potential benzenesulfonamide structures in each sample using pink color.
@@ A.4 Visualization of Top-50 Selection
A.5 Visualization on Selected Cases Containing Pyrimidine Sulfonamide

### v2

@@ A.2 Gradient Analysis
Appendix B Experimental Settings
B.1 Virtual Docking for Dataset Construction
We employed molecular docking to define the three-dimensional conformations of molecules within our DEL datasets. This method was applied to both the training and evaluation sets, generating ligand binding poses for all molecules. We concentrated on three pivotal carbonic anhydrase proteins: Q16790 (CAH9_HUMAN), P00918 (CAH2_HUMAN), and O43570 (CAH12_HUMAN).
For the Q16790 target, we sourced the 5fl4 and 2hkf PDB structures from the PDBbind database and utilized the Gerry dataset (Gerry et al., 2019). which comprised 108,529 molecules, generating up to nine potential poses per molecule. For the targets P00918 and O43570, we selected 127,500 SMILES strings from the DEL-MAP dataset (Hou et al., 2023) and conducted self-docking using the 3p3h and 5doh PDB structures for P00918, and 4kp5 and 4ht2 for O43570, as sourced from PDBbind. For the validation set, we applied the same docking methodology to the corresponding ligands of CA9, CA2, and CA12, involving 3,324, 6,395, and 2,690 ligands respectively.
In the specific docking procedures, initial 3D conformations of ligands were created using RDKit. The binding sites in the protein-ligand complexes were identified using 3D structural data of known binding ligands from PDBbind as reference points. Targeted docking was performed by defining the search space as a 22.5 Å cube centered on the reference ligand in the corresponding PDBbind complex. Using SMINA docking software, we generated 9 potential poses for each protein-ligand pair.
B.2 Hyperparameter Setting
@@
Hyper-parameter Selection
The hyperparameter configuration in the appendix requires clarification regarding the weight settings. The key parameters include $\mathcal{L}_{\text{rank}}$ weight, $\mathcal{L}_{\text{PSR}}$ weight, $\mathcal{L}_{\text{LGR}}$ weight, and ARC weight. The $\mathcal{L}_{\text{rank}}$ weight is logarithmically distributed between 1e9 and 1e11 to align with the magnitude of ZIP loss. $\mathcal{L}_{\text{PSR}}$ and $\mathcal{L}_{\text{LGR}}$ weights are calibrated to maintain appropriate balance among different ranking objectives. Given that ARC loss naturally aligns with ZIP loss magnitude, its weight is simply set to 1.0 or 0.1.
Temperature settings are determined by the characteristics of DEL read count data distribution, with denser distributions requiring lower temperatures. A detailed analysis of read count distribution and supporting theoretical proposition are provided in the Section C.1. Besides, the contract weight and margin serve as penalty terms for $\mathcal{L}_{\text{LGR}}$, with the weight selected based on $\mathcal{L}_{\text{LGR}}$’s relative magnitude. Detailed in Table 5, these values remain stable and consistent across experiments.
@@
$\mathcal{L}_{\text{consist}}$ weight $\gamma$ 1 0.1 0.1
$\mathcal{L}_{\text{rank}}$ weight $\rho$ 1.00e111.00E111.001011start_ARG 1.00 end_ARG start_ARG  end_ARG start_ARG e start_ARG 11 end_ARG end_ARG 1.00e91.00E91.00109start_ARG 1.00 end_ARG start_ARG  end_ARG start_ARG e start_ARG 9 end_ARG end_ARG 1.00e101.00E101.001010start_ARG 1.00 end_ARG start_ARG  end_ARG start_ARG e start_ARG 10 end_ARG end_ARG 1.00e81.00E81.00108start_ARG 1.00 end_ARG start_ARG  end_ARG start_ARG e start_ARG 8 end_ARG end_ARG 1.00e81.00E81.00108start_ARG 1.00 end_ARG start_ARG  end_ARG start_ARG e start_ARG 8 end_ARG end_ARG
$\mathcal{L}_{\text{PSR}}$ weight $\beta$ 0.5 0.91 0.91 0.67 0.5
$\mathcal{L}_{\text{LGR}}$ weight 1-$\beta$ 0.5 0.09 0.09 0.33 0.5
Temperature $T$ 0.8 0.3 0.2 0.9 0.2
$\mathcal{L}_{\text{con}}$ weight $\sigma$ 1.00e31.00E-31.0010-3start_ARG 1.00 end_ARG start_ARG  end_ARG start_ARG e start_ARG - 3 end_ARG end_ARG 1.00e31.00E-31.0010-3start_ARG 1.00 end_ARG start_ARG  end_ARG start_ARG e start_ARG - 3 end_ARG end_ARG 1.00e31.00E-31.0010-3start_ARG 1.00 end_ARG start_ARG  end_ARG start_ARG e start_ARG - 3 end_ARG end_ARG 1.00e41.00E-41.0010-4start_ARG 1.00 end_ARG start_ARG  end_ARG start_ARG e start_ARG - 4 end_ARG end_ARG 1.00e31.00E-31.0010-3start_ARG 1.00 end_ARG start_ARG  end_ARG start_ARG e start_ARG - 3 end_ARG end_ARG
Margin $\tau$ 1.00e31.00E-31.0010-3start_ARG 1.00 end_ARG start_ARG  end_ARG start_ARG e start_ARG - 3 end_ARG end_ARG 1.00e31.00E-31.0010-3start_ARG 1.00 end_ARG start_ARG  end_ARG start_ARG e start_ARG - 3 end_ARG end_ARG 1.00e31.00E-31.0010-3start_ARG 1.00 end_ARG start_ARG  end_ARG start_ARG e start_ARG - 3 end_ARG end_ARG 1.00e31.00E-31.0010-3start_ARG 1.00 end_ARG start_ARG  end_ARG start_ARG e start_ARG - 3 end_ARG end_ARG 1.00e31.00E-31.0010-3start_ARG 1.00 end_ARG start_ARG  end_ARG start_ARG e start_ARG - 3 end_ARG end_ARG
Proposition 1.
As $T\rightarrow 0$, the model simultaneously achieves: (1) Near-deterministic selection of compounds with the highest read counts, corresponding to the highest binding affinities; (2) Increased robustness to small noise perturbations in the DEL experiment data.
Based on the Proposition, the adaptive-ranking model would obtain more consistent identification of high-affinity compounds, reducing errors due to random fluctuations. Also, it achieves enhanced robustness against common DEL experimental noises such as PCR bias and sequencing errors. While lowering the temperature leads to a more deterministic ranking with high-affinity sensitivity and noise resistance, there exists overlooking of compounds with slightly lower rankings when the temperature goes to extremely low. In experiments, we demonstrate that [0.1, 0.4] should be a proper range for the distribution sharping.
Appendix C Experimental Results
C.1 Ablation Study on Hyper-parameters
In order to evaluate the robustness of our method, we conduct a comprehensive analysis of four critical hyperparameters: the consistency loss weight $\gamma$, ranking loss weight $\rho$, LGR loss weight $\beta$, and temperature $T$ across three datasets (3p3h, 4kp5-A, and 4kp5-OA). As shown in Table 6, we employ logarithmic search spaces for all loss-related hyperparameters to align the magnitudes of ranking and consistency losses with the ZIP loss, while adopting a linear search space for temperature.
The empirical results demonstrate that our selected hyperparameters consistently achieve optimal performance across all search spaces. The model exhibits strong stability, with performance variations remaining minimal under most hyperparameter adjustments. Nevertheless, we observe dataset-specific sensitivities: the 4kp5-OA dataset shows increased sensitivity to ranking loss weight variations, potentially due to elevated read count noise levels. Similarly, the 4kp5-A dataset exhibits performance fluctuations at higher values of ranking loss and $\mathcal{L}_{\text{LGR}}$ weights, which we attribute to magnitude imbalances in the numerical representations.
The performance progression with respect to temperature demonstrates a consistent linear relationship, providing empirical support for our distribution sharpening hypothesis. These findings collectively indicate that while our model maintains robustness across the hyperparameter search space with well-justified parameter selections, its sensitivity can be influenced by dataset-specific characteristics, particularly read count distribution noise and magnitude disparities in the underlying data.
Table 6: Comparison of different hyper-parameters on binding affinity prediction performance. The best performance within one set of hyperparameter group is set bold.
Parameter Value 3p3h (CA2) 4kp5-A (CA12) 4kp5-OA (CA12)
Metric Sp SubSp Sp SubSp Sp SubSp
0.1 -0.275±0.011 -0.163±0.017 -0.268±0.012 -0.277±0.016 -0.289±0.025 -0.233±0.021
$\mathcal{L}_{\text{consist}}$ weight $\gamma$ 1 -0.286±0.002 -0.177±0.005 -0.266±0.008 -0.238±0.008 -0.287±0.005 -0.213±0.014
10 -0.276±0.010 -0.163±0.015 -0.258±0.019 -0.239±0.010 -0.278±0.024 -0.227±0.040
1.00e091.00E+091.001009start_ARG 1.00 end_ARG start_ARG  end_ARG start_ARG e start_ARG 09 end_ARG end_ARG -0.266±0.011 -0.151±0.016 -0.268±0.012 -0.277±0.016 -0.152±0.045 -0.225±0.023
$\mathcal{L}_{\text{rank}}$ weight $\rho$ 1.00e101.00E+101.001010start_ARG 1.00 end_ARG start_ARG  end_ARG start_ARG e start_ARG 10 end_ARG end_ARG -0.269±0.006 -0.151±0.009 -0.257±0.005 -0.189±0.016 -0.289±0.025 -0.233±0.021
1.00e111.00E+111.001011start_ARG 1.00 end_ARG start_ARG  end_ARG start_ARG e start_ARG 11 end_ARG end_ARG -0.286±0.002 -0.177±0.005 -0.135±0.012 -0.060±0.036 -0.084±0.095 -0.058±0.077
0.09 -0.277±0.009 -0.165±0.013 -0.268±0.012 -0.277±0.016 -0.289±0.025 -0.233±0.021
$\mathcal{L}_{\text{LGR}}$ weight $\beta$ 0.5 -0.286±0.002 -0.177±0.005 -0.267±0.033 -0.240±0.016 -0.288±0.025 -0.247±0.019
0.91 -0.275±0.011 -0.160±0.019 -0.173±0.054 -0.089±0.038 -0.279±0.007 -0.222±0.033
0.2 -0.280±0.021 -0.173±0.029 -0.267±0.013 -0.247±0.009 -0.289±0.025 -0.233±0.021
Temperature $T$ 0.5 -0.279±0.009 -0.169±0.014 -0.266±0.014 -0.236±0.012 -0.275±0.013 -0.216±0.005
0.8 -0.286±0.002 -0.177±0.005 -0.268±0.010 -0.222±0.010 -0.275±0.035 -0.220±0.029
1 -0.279±0.011 -0.166±0.015 -0.247±0.022 -0.265±0.014 -0.256±0.033 -0.181±0.046
C.2 Ablation Study on Structure Information
To assess the value of structural information from docking software and its complementarity with sequence features, we performed an ablation study focusing on the additive combination of structure and fingerprint embeddings in the ARC algorithm. We applied varying scaling factors (0, 0.3, 0.6, 1.0, 1.5, and 2.0) to the structure embedding across three datasets (3p3h, 4kp5-A, and 4kp5-OA) with five random seeds. Table 7 shows that incorporating structural information significantly improves model performance. The analysis revealed higher model sensitivity in the noise-prone 4kp5-OA dataset, while performance degradation was observed in 4kp5-A when scaling factors exceeded 1.0. These results indicate that while structural information enhances model performance, excessive weighting of potentially uncertain structural data can impair predictions. Nevertheless, our chosen parameterization demonstrates consistent performance across all datasets.
Table 7: Parameter value comparison for structure scaling factor. The best performance within one set of hyperparameter group is set bold.
Value $\varsigma$ 3p3h (CA2) 4kp5-A (CA12) 4kp5-OA (CA12)
Metric Sp SubSp Sp SubSp Sp SubSp
0 -0.236±0.010 -0.112±0.013 -0.253±0.012 -0.218±0.017 -0.195±0.044 -0.103±0.055
0.3 -0.262±0.008 -0.145±0.012 -0.265±0.017 -0.227±0.017 -0.124±0.146 -0.062±0.090
0.6 -0.263±0.008 -0.146±0.011 -0.250±0.017 -0.231±0.019 -0.210±0.040 -0.121±0.047
1 -0.286±0.002 -0.177±0.005 -0.268±0.012 -0.277±0.016 -0.289±0.025 -0.233±0.021
1.5 -0.270±0.011 -0.155±0.016 -0.244±0.022 -0.$2$0.152±0.156 -0.139±0.104
2 -0.271±0.012 -0.155±0.015 -0.191±0.089 -0.216±0.060 -0.230±0.038 -0.152±0.051
Figure 4: Quantitative analysis of Top-50 selection, including $K_{i}$ distribution and accuracy for DEL-Dock (Shmilovich et al., 2023).
C.3 Comparison Result of Top-50 Selection Cases by DEL-Dock
To evaluate the DEL-Dock model’s performance, we analyzed the $K_{i}$ value distributions and benzenesulfonamide identification rates for the top-50 compounds across multiple datasets. Figure 4 demonstrates that DEL-Dock outperformed DEL-Ranking in the 3p3h dataset and showed comparable results in 4kp5-A and 5fl4-9p datasets, while DEL-Ranking exhibited superior performance in 4kp5-OA and 5fl4-20p datasets. The benzenesulfonamide identification accuracy mirrored these trends across all datasets, with DEL-Dock showing strength in 3p3h, equivalent performance in 4kp5-A and 5fl4-9p, and DEL-Ranking maintaining advantage in 4kp5-OA and 5fl4-20p datasets. Further analysis indicated that ranking-based methods performed better in datasets with higher noise levels and increased read counts, aligning with theoretical expectations.
Figure 5: Visualization of Top-50 high affinity cases without benzene sulfonamide.
Moreover, analysis of high-affinity compounds devoid of benzenesulfonamide functional groups was performed across the 4kp5-OA, 5fl4-9p, and 5fl4-20p datasets. The identification of thiocarbonyl and sulfonamide groups exhibiting Ki values below 10.0 yielded two significant insights. First, it validated the benzenesulfonamide-based activity labeling approach, particularly given that DEL-Dock achieved these results independent of activity label data. Second, the label-guided DEL-Ranking model successfully identified structurally analogous functional groups to benzenesulfonamides, demonstrating that ranking supervision effectively enhances the discovery of novel, high-activity molecular scaffolds.
C.4 Visualization of Top-50 Selection of DEL-Ranking
@@
C.5 Visualization on Selected Cases Containing Pyrimidine Sulfonamide